OpenStudy (anonymous):

X^2 -20x + 100 = ? Can some help me please?

OpenStudy (anonymous):

\[x^2-20x+100=(x-10)(x-10)=(x-10)^{2}\]

OpenStudy (anonymous):

Thank You :)

OpenStudy (anonymous):

You welcome !